# Exploration and Anti-Exploration with Distributional Random Network Distillation

### v1

@@ Abstract
Exploration remains a critical issue in deep reinforcement learning for an agent to attain high returns in unknown environments. Although the prevailing exploration Random Network Distillation (RND) algorithm has been demonstrated to be effective in numerous environments, it often needs more discriminative power in bonus allocation. This paper highlights the “bonus inconsistency” issue within RND, pinpointing its primary limitation. To address this issue, we introduce the Distributional RND (DRND), a derivative of the RND. DRND enhances the exploration process by distilling a distribution of random networks and implicitly incorporating pseudo counts to improve the precision of bonus allocation. This refinement encourages agents to engage in more extensive exploration. Our method effectively mitigates the inconsistency issue without introducing significant computational overhead. Both theoretical analysis and experimental results demonstrate the superiority of our approach over the original RND algorithm. Our method excels in challenging online exploration scenarios and effectively serves as an anti-exploration mechanism in D4RL offline tasks.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Bonus Heatmap of dataset distribution and RND bonus. The left image illustrates the dataset distribution, the middle image represents the RND bonus before training, and the right image represents the RND bonus after training. A more detailed change process is in Appendix E.8. Ideally, we aim for a uniform bonus distribution before any training and without exposure to the dataset. After extensive training, the expected bonus should inversely correlate with the dataset distribution. The bonus distribution of RND is inconsistent with the desired distribution, indicating a problem with bonus inconsistency. The experiment details can be found in Appendix C
@@ 3 Preliminaries
Here $\lambda$ is the scale of the bonus for the update of the value net.
@@ 4.2 Distill the target network of random distribution
Unlike RND, which only has one target network $\bar{f}(s)$, the DRND algorithm has $N$ target networks $\bar{f}_{1}(s,a),\bar{f}_{2}(s,a),...,\bar{f}_{N}(s,a)$, which are from a random distribution with randomly initialized parameters and do not participate in training. In DRND, we use $s$ as input in the online setting and $(s,a)$ pair as input in the offline setting. For simplicity, we define $x=(s,a)$ (offline setting) or $x=(s)$ (online setting). For each state-action pair $x$, we construct a variable $c(x)$ that satisfies the distribution:
@@ 4.4 Bonus of the DRND agent
In summary, the total bonus, as seen in Equations 6 and 9, is
$b(s_{t},a_{t})=\alpha\|f_{\theta}(x)-\mu(x)\|^{2}+(1-\alpha)\sqrt{\frac{[f_{\theta}(x)]^{2}-[\mu(x)]^{2}}{B_{2}(x)-[\mu(x)]^{2}}}.$ (10)
@@
where $\alpha$ represents the scaling factor for the two bonus terms. Figure 2 shows the diagram of our method and the RND method. For smaller values of $n$, the variance of the second bonus estimate is substantial, rendering the first bonus a more dependable measure for states with infrequent occurrences. Conversely, as $n$ increases, the variance of the second bonus approaches zero, enhancing its reliability. It’s worth noting that during DRND predictor updates, the first bonus diminishes quicker than the second. This dynamic permits a constant $\alpha$ to attain the intended behaviour, as illustrated in Figure 3.
The DRND’s target network remains static, and its loss and intrinsic reward calculations do not introduce new backpropagation, keeping computational time similar to the RND algorithm. With $\alpha=1$ and $N=1$, the loss and intrinsic reward simplify to $\left\|f_{\theta}(x)-\bar{f}(x)\right\|^{2},$ aligning with RND. Unlike count-based and pseudo-count methods, we do not use an extra network or table for state occurrences but estimate using the prediction network’s statistics. When $\alpha=0$ and $c(x)\sim\begin{array}[]{c|c c }\hline\cr X&-1&1\\ \hline\cr P&0.5&0.5\\ \hline\cr\end{array}$, the loss and intrinsic reward simplify to $\|f_{\theta}(x)\|^{2}$ and $\|f_{\theta}(x)\|$, aligning with the pseudo-count approach CFN.
@@ Appendix C Implementation Details and Experimental Settings
The architecture of predictor networks in Figure 1 and Figure 3 is 3 linear layers, the input dim is 2, hidden dim and output dim is 16, activate function is ReLU. Target networks’ architecture is 2 linear layers, the input dim is 2, hidden dim and output dim is 16, activate function is ReLU.
@@ E.3 Expected Online Performance
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ E.4.1 Online Tasks
We conduct the adversarial attack experiments with different numbers of target networks in DRND. As shown in Figure 9, the robustness of DRND generally improves with an increase in the target number $N$. Considering both runtime and performance, we chose $N=10$ as the optimal number of targets for our online experiments.
@@ E.4.2 Offline Tasks
The results are shown in Table 8. The results indicate that the average score demonstrates an upward trend as the number of targets increases. At the same time, its variance decreases, which suggests that a higher number of targets generally leads to improved and more consistent outcomes. However, it’s worth noting that there are diminishing returns; for instance, the differences between the results at $N=10$ and $N=20$ are marginal. Considering these considerations, we chose $N=10$ for our offline experiments. Furthermore, the algorithm exhibits limited sensitivity to variations in the number of targets in online and offline settings.
@@ E.6.2 Offline Tasks
We examine the influence of $\alpha$ on offline tasks using the D4RL dataset. We employ various values of $\alpha$ to train an offline agent on the ‘medium’ datasets. The final scores are presented in Table 9, and the training curves are shown in Figure 12. It is observed that in some cases, when $\alpha=0.9$, the final score is higher, and the training curve exhibits greater stability. Consequently, we consistently opted for $\alpha=0.9$ in our offline experiments. When $\alpha=1$, only the first bonus term comes into play, and the results are not as favorable as when $\alpha=0.9$, demonstrating the effectiveness of the second bonus term. Additionally, when examining the final results, it becomes evident that our first bonus outperforms the RND.
@@
E.7 Evaluation on offline-to-online D4RL
@@
E.8 More Detailed Change Process of RND Bonus

### v4

@@ Abstract
Exploration remains a critical issue in deep reinforcement learning for an agent to attain high returns in unknown environments. Although the prevailing exploration Random Network Distillation (RND) algorithm has been demonstrated to be effective in numerous environments, it often needs more discriminative power in bonus allocation. This paper highlights the “bonus inconsistency” issue within RND, pinpointing its primary limitation. To address this issue, we introduce the Distributional RND (DRND), a derivative of the RND. DRND enhances the exploration process by distilling a distribution of random networks and implicitly incorporating pseudo counts to improve the precision of bonus allocation. This refinement encourages agents to engage in more extensive exploration. Our method effectively mitigates the inconsistency issue without introducing significant computational overhead. Both theoretical analysis and experimental results demonstrate the superiority of our approach over the original RND algorithm. Our method excels in challenging online exploration scenarios and effectively serves as an anti-exploration mechanism in D4RL offline tasks. Our code is publicly available at https://github.com/yk7333/DRND.
Reinforcement Learning, Exploration, Anti-exploration
@@ 1 Introduction
Figure 1: Bonus Heatmap of dataset distribution and RND bonus. The left image illustrates the dataset distribution, the middle image represents the RND bonus before training, and the right image represents the RND bonus after training. A more detailed change process is in Appendix E.9. Ideally, we aim for a uniform bonus distribution before any training and without exposure to the dataset. After extensive training, the expected bonus should inversely correlate with the dataset distribution. The bonus distribution of RND is inconsistent with the desired distribution, indicating a problem with bonus inconsistency. The details of the experiment settings can be found in Appendix C
@@ 3 Preliminaries
where $\lambda$ is the scale of the bonus for the update of the value network.
@@ 4.2 Distill the target network of random distribution
Unlike RND, which only has one target network $\bar{f}$, the DRND algorithm has $N$ target networks $\bar{f}_{1},\bar{f}_{2},...,\bar{f}_{N}$, which are from a random distribution with randomly initialized parameters and do not participate in training. In DRND, we use $s$ as input in the online setting and $(s,a)$ pair as input in the offline setting. For simplicity, we define $x=(s,a)$ (offline setting) or $x=(s)$ (online setting). For each state-action pair $x$, we construct a variable $c(x)$ that satisfies the distribution:
@@
Table 1: Comparison of the bonus distributions of RND and DRND against (a) the uniform distribution $U$, (b) the distribution of $1/\sqrt{n}$. $P$ denotes the distribution of RND bonus or DRND bonus.
Metrics RND DRND
$D_{KL}(P\|U)$ (Before training) 0.0377$\pm$0.0248 0.0070$\pm$0.0063
$D_{KL}(P\|1/\sqrt{n})$ (After training) 0.0946$\pm$0.0409 0.0476$\pm$0.0389
Table 2: KL-divergence comparison of the distributions of overall bonus $b$, individual bonus terms $b_{1}$, $b_{2}$ against (a) the uniform distribution $U$, (b) the distribution of $1/\sqrt{n}$. $P$ represents the bonus distribution of $b$, $b_{1}$ or $b_{2}$.
Metrics $b$ $b_{1}$ $b_{2}$
$D_{KL}(P\|U)$ (Before training) 0.0067$\pm$0.0035 0.0070$\pm$0.0038 0.0104$\pm$0.0055
$D_{KL}(P\|1/\sqrt{n})$ (After training) 0.0524$\pm$0.0248 0.0703$\pm$0.0404 0.0396$\pm$0.0209
@@ 4.4 Bonus of the DRND agent
The total bonus that combines Equations 6 and 9 gives
$b(x)=\alpha\|f_{\theta}(x)-\mu(x)\|^{2}+(1-\alpha)\sqrt{\frac{[f_{\theta}(x)]^{2}-[\mu(x)]^{2}}{B_{2}(x)-[\mu(x)]^{2}}},$ (10)
@@
where $\alpha$ represents the scaling factor for the two bonus terms. Figure 2 shows the diagram of our method and the RND method. For smaller values of $n$, the variance of the second bonus estimate is substantial, rendering the first bonus $b_{1}$ a more dependable measure for states with infrequent occurrences. Conversely, as $n$ increases, the variance of the second bonus $b_{2}$ approaches zero, enhancing its reliability. From the experiments, we found that during the initial training phase, the magnitudes of the two bonuses, denoted as $b_{1}$ and $b_{2}$, are nearly the same. However, as training progresses, the rate at which the first bonus $b_{1}$ decreases is much faster than that of the second bonus $b_{2}$. Eventually, the magnitude of $b_{1}$ becomes approximately two orders of magnitude lower than $b_{2}$. Therefore, if we want $b_{1}$ to dominate during the early stages of training and have $b_{2}$ dominate during the later stages, we do not need to set a dynamic coefficient to achieve this functionality. Instead, we can simply set it as a fixed constant $\alpha$. The distribution of various bonuses as well as the total bonus of the DRND method before and after training can be seen in Figure 3.
The aforementioned heatmap serves as a qualitative illustration. We now provide quantitative experimental results to demonstrate that DRND offers superior bonuses compared to RND. In count-based exploration methods, it is common to assess the rationality of the bonus distribution by comparing it with the $1/\sqrt{n}$ distribution, where $n$ represents the state visitation count. We evaluate the discrepancy between the intrinsic reward distribution and the uniform distribution $U$ before training, and the distribution of $1/\sqrt{n}$ after training using KL divergence. We randomly select 100 initial dataset distributions and report the mean and variance of the KL divergence. This assessment is denoted as $D_{KL}(P\|U)$ and $D_{KL}(P\|1/\sqrt{n})$, where $P$ represents the distribution of RND bonus or DRND bonus. The results obtained are presented in Table 1. It is evident that the KL divergence between the bonuses of DRND and the uniform distribution is smaller during the initial training stages, effectively promoting uniform exploration by the agent. In the later stages of training, DRND better aligns with the distribution of $1/\sqrt{n}$, thereby facilitating deeper exploration by the agent.
To validate that $b_{1}$ gives better and more uniform bonuses during the early stages of training, while $b_{2}$ better captures the state visitation count $n$ later on, we similarly measure the KL divergence between each bonus term $b_{1}$, $b_{2}$, the overall bonus $b$ and (a) the uniform distribution $U$ before training; (b) the distribution of $1/\sqrt{n}$ after training. The results are presented in Table 2.
From the table, it can be observed that in the early stages of training, $b_{1}$ performs better, providing a better solution to the initial bonus inconsistency. Conversely, in the later stages of training, $b_{2}$ performs better, addressing the final bonus inconsistency more effectively. Moreover, from the KL divergence of the total bonus, it is evident that setting a fixed value for $\alpha$ achieves the desired effect of having $b_{1}$ dominate during the early stages and $b_{2}$ dominate during the later stages, without the need for dynamic adjustments.
The combination of $b_{1}$ and $b_{2}$ depends on the environment and task at hand. In some environments, $b_{1}$ may be more crucial, while in others, the contribution of $b_{2}$ might be greater. However, regardless of the specifics, properly integrating both can lead to improved performance.
4.5 Connections between DRND and prior methods
The target network of the DRND method remains static throughout the training process. Moreover, its loss and intrinsic reward computations do not entail additional backpropagation steps, thereby maintaining computational efficiency comparable to that of the RND algorithm. Specifically, when the hyperparameters $\alpha$ and $N$ are set to 1, the expressions for the loss and intrinsic reward become simplified to $\left\|f_{\theta}(x)-\bar{f}(x)\right\|^{2}$, which coincides with the formulation utilized in the RND framework. In contrast to count-based and pseudo-count methods, we forgo the utilization of an additional network or table to track state occurrences. Instead, we estimate these occurrences using the statistical information derived from the prediction network itself. When $\alpha$ is set to 0 and $c(x)$ follows the distribution $c(x)\sim\begin{array}[]{c|c c }\hline\cr X&-1&1\\ \hline\cr P&0.5&0.5\\ \hline\cr\end{array}$, the expressions for the loss and intrinsic reward are simplified to $\|f_{\theta}(x)\|^{2}$ and $\|f_{\theta}(x)\|$, which align with the pseudo-count approach CFN.
@@ 6 Conclusion
Acknowledgements
This work was supported by the STI 2030-Major Projects under Grant 2021ZD0201404.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Implementation Details and Experimental Settings
The architecture of predictor networks in Figure 1 and Figure 3 is 3 linear layers, the input dim is 2, hidden dim and output dim is 16, activate function is ReLU. Target networks’ architecture is 2 linear layers, the input dim is 2, hidden dim and output dim is 16, activate function is ReLU. We set the state as 2-dimensional, with the state space defined as the 2-dimensional space [0,1]x[0,1]. The given distribution is regarded as the dataset distribution of the current agent passing through states, and the dimensions can be arbitrary. Setting it as 2-dimensional is merely for visualization convenience.
@@ E.4.1 Online Tasks
We conduct adversarial attack experiments with different numbers of target networks in DRND. As shown in Figure 9, the robustness of DRND generally improves with an increase in the target number $N$. Considering both runtime and performance, we chose $N=10$ as the optimal number of targets for our online experiments.
@@ E.4.2 Offline Tasks
The results are shown in Table LABEL:target. The results indicate that the average score demonstrates an upward trend as the number of targets increases. At the same time, its variance decreases, which suggests that a higher number of targets generally leads to improved and more consistent outcomes. However, it’s worth noting that there are diminishing returns; for instance, the differences between the results at $N=10$ and $N=20$ are marginal. Considering these considerations, we chose $N=10$ for our offline experiments. Furthermore, the algorithm exhibits limited sensitivity to variations in the number of targets in online and offline settings.
@@ E.6.2 Offline Tasks
We examine the influence of $\alpha$ on offline tasks using the D4RL dataset. We employ various values of $\alpha$ to train an offline agent on the ‘medium’ datasets. The final scores are presented in Table LABEL:alpha, and the training curves are shown in Figure 12. It is observed that in some cases, when $\alpha=0.9$, the final score is higher, and the training curve exhibits greater stability. Consequently, we consistently opted for $\alpha=0.9$ in our offline experiments. When $\alpha=1$, only the first bonus term comes into play, and the results are not as favorable as when $\alpha=0.9$, demonstrating the effectiveness of the second bonus term. Additionally, when examining the final results, it becomes evident that our first bonus outperforms the RND.
@@
E.7 Parameter Study on $\lambda$
We present the results of our parameter study on $\lambda$, which are detailed in the table below. We choose some environments and show the results. All experiments are conducted across 5 seeds, and we report the mean and standard deviation of their results.
Table 12: Hopper-Medium-v2
$\lambda_{actor}$ $\lambda_{critic}$ $10$ $15$ $20$
$10$ $98.7\pm 2.0$
$15$ $95.9\pm 3.4$ $98.5\pm 1.1$ $96.1\pm 3.9$
$20$ $94.65\pm 2.6$
Table 13: Hopper-Medium-Replay-v2
$\lambda_{actor}$ $\lambda_{critic}$ $5$ $10$ $15$
$3$ $99.9\pm 1.3$
$5$ $94.7\pm 5.9$ $100.5\pm 1.0$ $99.3\pm 0.4$
$8$ $99.8\pm 1.2$
Table 14: HalfCheetah-Medium-v2
$\lambda_{actor}$ $\lambda_{critic}$ $0.05$ $0.1$ $0.2$
$0.5$ $65.4\pm 0.8$
$1$ $67.8\pm 0.4$ $68.3\pm 0.2$ $68.2\pm 1.1$
$2$ $68.0\pm 0.3$
Table 15: HalfCheetah-Full-Replay-v2
$\lambda_{actor}$ $\lambda_{critic}$ $0.5$ $1$ $2$
$0.5$ $81.2\pm 0.5$
$1$ $81.5\pm 0.7$ $81.4\pm 1.7$ $81.7\pm 1.3$
$2$ $80.9\pm 0.9$
Table 16: Walker-Full-Replay-v2
$\lambda_{actor}$ $\lambda_{critic}$ $1$ $3$ $5$
$1$ $107.2\pm 1.2$
$3$ $61.3\pm 25.8$ $109.6\pm 0.7$ $109.4\pm 0.4$
$5$ $109.4\pm 0.6$
Table 17: Walker-Medium-Expert-v2
$\lambda_{actor}$ $\lambda_{critic}$ $15$ $20$ $25$
$10$ $100.5\pm 5.6$
$15$ $93.9\pm 8.8$ $109.6\pm 1.0$ $85.2\pm 15.0$
$20$ $110.3\pm 0.6$
E.8 Evaluation on offline-to-online D4RL
@@
E.9 More Detailed Change Process of RND Bonus